# Characterizing GPU Resilience and Impact on AI/HPC Systems

### v3

@@ Abstract.
This study characterizes GPU resilience in Delta  [^1] Delta is an HPC system operated by the National Center for Supercomputing Applications (NCSA) at the University of Illinois Urbana-Champaign., a large-scale AI system that consists of 1,056 A100 and H100 GPUs, with over 1,300 petaflops of peak throughput. We used 2.5 years of operational data (11.7 million GPU hours) on GPU errors. Our major findings include: (i) H100 GPU memory resilience is worse than A100 GPU memory, with 3.2x lower per-GPU MTBE for memory errors, (ii) The GPU memory error-recovery mechanisms on H100 GPUs are insufficient to handle the increased memory capacity, (iii) H100 GPUs demonstrate significantly improved GPU hardware resilience over A100 GPUs with respect to critical hardware components, (iv) GPU errors on both A100 and H100 GPUs frequently result in job failures due to the lack of robust recovery mechanisms at the application level, and (v) We project the impact of GPU node availability on larger-scales and find that significant overprovisioning of 5% is necessary to handle GPU failures.
@@
*Equal Contribution
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 000-0-$0$-$0$-0/00/00
@@ 1. Introduction
Large-scale HPC systems are important not only for scientific workloads (Skinner and Kramer, 2005) but also for data analytics (Alam et al., 2022) and machine learning (ML) (Huyen, 2022). The main components of these systems are specialized accelerators, such as GPUs, that enable acceleration of computations, such as ML training (Touvron et al., 2023; Anil et al., 2023), ML inference (Li et al., 2024), and simulations (Phillips et al., 2005; Shainer et al., 2009).
This paper studies the resilience of A100 GPUs and compares the result with H100 GPUs of the GH200 Grace Hopper Superchips [^2] GH200 Superchip, hereafter., together with their associated memory: 40 GB HBM2e on each A100 GPU, and 96 GB HBM3 on each H100 GPU, respectively. The A100 GPU nodes and H100 GPUs (GH200 Superchip) nodes are operated as two independent systems sharing a storage cluster running the Lustre file system, allowing us to study and compare the two systems (refer to as Delta). The workflow on Delta involves users from universities nationwide and presents a spectrum of HPC and ML workloads. The study uses 2.5 years of data on critical GPU errors collected across the stated GPUs, encompassing 9.6 million GPU hours of A100 GPUs and 2.1 million GPU hours of H100, a combined 11.7 million GPU hours.
@@
Figure 1. A double-bit memory error (XID 48) occurred, and it is uncorrectable by the SECDED ECC HBM3 memory. Due to this double-bit error, the user job scheduled on that GPU failed, as reflected in the scheduler logs. Subsequently, this uncorrectable memory error requires a node draining and reset to complete the row remapping recovery action. The total recovery process for this incident took 19 hours, during which the node was unavailable for accepting new jobs. This incident shows that a GPU error can lead to user job failures and significantly impact node availability.
[…]
A code snippet showing GSP Error Caused App Failure.
@@
    H100 shows 3.2×3.2$\times$3.2 × lower per-GPU mean time between errors (MTBE) compared to A100 for uncorrectable ECC memory errors. The per-GB MTBE of the H100’s HBM3 memory is 24% lower ($\sim\!8.5$M hours) than the A100’s HBM2e memory ($\sim\!11.3$M hours). We conjecture that the reduction in memory resilience stems from H100’s higher memory capacity.
  2. (ii)
@@
    The GPU memory error-recovery mechanisms on A100 and H100 GPUs (e.g., memory row remapping, error containment) (NVIDIA Corporation, 2024) improve GPU memory resilience and reduce service interruption. We observed that these mechanisms mitigate (e.g., using memory row remapping) 92% of uncorrectable ECC memory errors on H100 GPUs. However, the memory error-recovery mechanism is insufficient to handle the increase in memory capacity and the corresponding increase in row remapping events on the H100 GPUs.
  3. (iii)
@@
    H100 GPUs demonstrate significantly improved hardware resilience over A100 GPUs, with respect to critical components such as GSP [^3] A GPU system processor (GSP) is an onboard co-processor that offloads driver tasks from the CPU for latency and performance improvement., NVLink, and PMU SPI [^4] PMU on an NVIDIA’s GPU regulates the frequency, voltage, and power of the GPU based on various factors such as temperature and power cap. SPI stands for serial peripheral interface, which serves as the communication channel between peripheral hardware., which were major sources of job failures in A100 systems. We attribute this to driver-level enhancements and tighter integration (NVIDIA, 2023; NVIDIA Corporation, 2023), which contribute to improved resilience. Specifically, comparing H100 and A100 GPU hardware, we observed (a) a significant reduction of GSP errors on H100 (only 3 cases in our measurement period) and (b) the elimination of PMU SPI error propagations, which on A100 GPUs can lead to MMU errors 88% of times with 90% leading to user job failures, (c) no NVLink errors on H100 GPUs during the measurement period.
  4. (iv)
@@
    GPU errors on both A100 and H100 GPUs frequently result in job failures due to the lack of robust recovery mechanisms at the application level. Except for MMU and NVLink errors, other GPU errors cannot be handled by application-level mechanisms, resulting in close to $100\%$ job failure rate. The underlying cause of job failures differs by GPU type: hardware errors are the predominant cause in A100 GPUs, whereas memory errors are the primary cause in H100 GPUs. Overall, we find that application-based recovery strategies are largely ineffective; hence, there is a compelling need to improve resilience at the GPU memory and hardware level.
  5. (v)
@@
    The overall availability per-GPU node is approximately $99.4\%$ for A100 GPUs and $99.3\%$ for H100 GPUs, corresponding to a downtime between 9–10 minutes per day. We projected the impact of this measured availability on larger scales via emulation. For example, to maintain 99.9% availability at the job level, overprovisioning of 5% would be necessary. While at first glance, such overprovisioning would appear to be a small cost, for the above example, it would cost over \$1 million per month. If GPU node availability were improved to 99.9%, the required overprovisioning would reduce by 2.5×2.5$\times$2.5 ×.
@@
Putting the paper in perspective. Previous studies on characterizing GPU resilience in large-scale systems (DeBardeleben et al., 2014; Di Martino et al., 2014; Tiwari et al., 2015b, a; Gupta et al., 2015, 2017; Ostrouchov et al., 2020; Nie et al., 2016, 2017; Oles et al., 2024) focus on GPU memory errors in older GPU generations (Tesla, Kepler, and Volta) that lack the latest resilience (e.g., row remapping, error containment, NVLink CRC-retry) and performance features (e.g., GPU System Processor) introduced in NVIDIA Ampere-generation GPUs. A recent study from Meta (Kokolis et al., 2025) characterizes cluster-level resilience for two large-scale machine learning clusters equipped with A100 GPUs. Our paper provides a deeper understanding of GPU errors and failures of two recent generations of GPUs, A100 and H100, and their impact on a broad set of HPC/ML applications. To the best of our knowledge, this is the first study on A100 and H100 GPU errors in HPC/ML systems.
@@ 2. Background
This section provides information on (i) Delta specification and overall GPU utilization, (ii) critical GPU error categories used in this study, and (iii) GPU error management and recovery.
Figure 2. System architecture and specifications of Delta. This study focuses on the H100 and A100 GPU nodes.
[…]
Delta System Diagram.
2.1. Delta Specifications and Utilization
@@
Delta GPU Utilization. Delta’s NVIDIA A100 GPUs are frequently scheduled and utilized, with an average GPU utilization of 51% during the operational period. NVIDIA H100 GPUs show a slightly lower average utilization (41%) than A100 GPUs.
Figure 3. NVIDIA memory error recovery process for A100 and H100 GPUs.
[…]
A plot showing NVIDIA GPU recovery path
@@ 2.2. NVIDIA GPU Error Categories
NVIDIA GPU errors are reported as XID errors. In this study, we selected a subset of XID errors that are described as common and high-impact by NVIDIA’s Developer Manuals (NVIDIA, 2024; NVIDIA Corporation, 2024), NVIDIA Developer Forums and Blogs, and Delta site reliability engineers (SREs). We primarily collected errors and their associated recovery events. The selected XID errors/events indicate GPU issues that often cannot be resolved without SRE’s interventions (e.g., node service and GPU replacement). The selected XIDs and their corresponding GPU errors are described in Section 4,  Table 1. We categorize the selected GPU errors into three categories: (i) GPU hardware, which includes all onboard hardware except for GPU memory and NVLink interconnect, (ii) NVLink interconnect, and (iii) GPU memory. Note that General GPU Software Error (XID 13) and Reset Channel Verification Error (XID 43) are usually caused by user jobs and do not impact the health of the GPU (NVIDIA, 2024); we excluded those errors from our study.
GPU Hardware Errors. The critical GPU hardware errors we studied include MMU [^5] The memory management unit (MMU) provides essential memory I/O functionalities. errors, GPU Fallen Off the Bus errors, GSP errors, and PMU communication errors. We do not consider other GPU hardware errors in our study. GPU hardware errors can lead to user job failures, GPU halt, and data corruption. Among those errors, GPU Fallen Off the Bus and GSP errors lead to GPU failures, and manual GPU resets or node reboots are required to recover from the error (NVIDIA, 2024). Delta SREs monitor GSP errors closely to ensure timely recovery to maintain GPU availability.
GPU Interconnect (NVLink) Errors. GPU-GPU NVLink errors are caused by faulty GPU hardware, connectors, or improper connector installation during system integration, and can lead to GPU unavailability and user job failures. NVLink errors impede data transfer between GPUs and reduce computational throughput. A GPU reset or node reboot is required to clear NVLink errors (NVIDIA, 2024).
GPU Memory Errors. GPU memory errors included in this study are double-bit errors (DBEs) and consecutive single-bit errors (SBEs) [^6] Consecutive SBEs are multiple single-bit error (SBE) occurrences at the same memory location.. Individual SBEs are not logged, as they are automatically corrected by ECC. DBEs and consecutive SBEs are considered uncorrectable ECC memory errors by the NVIDIA driver, and they trigger downstream error-recovery mechanisms (NVIDIA, 2024; NVIDIA Corporation, 2024), which are introduced in Section 2.1. Failures in these mechanisms can lead to GPU/node failures and require GPU or node reboots to recover (NVIDIA, 2024). Delta SREs continuously monitor uncorrectable ECC memory errors and error-recovery failures to ensure timely replacement of faulty GPUs.
@@ 2.3. NVIDIA GPU Error Management
GPU Memory. Figure 3 shows the uncorrectable ECC memory error-recovery process (NVIDIA Corporation, 2024) for A100 and H100 in more detail. The primary mechanism for mitigating uncorrectable ECC memory errors for A100 and H100 GPUs is row-remapping, wherein the faulty memory row is replaced with a spare row, and a row-remapping event (RRE) is logged. The actual row remapping happens at the next GPU reset (e.g., during node reboot or maintenance). If there are no spare memory rows, a row remapping failure (RRF) is indicated  (NVIDIA, 2024; NVIDIA Corporation, 2024).
A100 and H100 GPUs support online recovery mechanisms such as error containment and dynamic page offlining (NVIDIA, 2024; NVIDIA Corporation, 2024) for mitigating uncorrectable ECC memory errors with minimal node interruption. The dynamic page offlining marks the faulty memory page as unusable without requiring a GPU reset to maintain availability. The error containment procedure terminates user processes using the faulty memory address to prevent error propagation to other applications. Successful error containment is logged as a Contained Memory Error, whereas an unsuccessful error containment is logged as an Uncontained Memory Error. Failure in a row-remapping or error containment can cause a GPU failure that requires a GPU reset or node reboot. Delta SREs monitor row-remapping failures and replace GPUs that repeatedly emit such errors.
@@ 3.1. Data Sources
Figure 4. Overview of our data collection, processing, and analysis pipeline.
[…]
Pipeline figure showing data collection, processing and analysis.
@@
System logs. System logs collected from all compute nodes capture events across system components. We created a set of regular expression (RegEX) patterns and used it to extract GPU error-recovery log entries by referring to NVIDIA XID messages (NVIDIA, 2024) from the system logs (Figure 4, item (1)). The GPU error logs were our major sources of error and recovery information.
Slurm scheduler database. Delta uses the Slurm Workload Manager (Yoo et al., 2003) (“Slurm scheduler” hereafter) for scheduling user jobs. The Slurm scheduler database keeps track of user job scheduling events, including the start and end times, the scheduled nodes, resource usage, job status, exit codes, and the srun command line. We used the Slurm database for user job failure characterization.

### v4

# Story of Two GPUs: Characterizing the Resilience of Hopper H100 and Ampere A100 GPUs

@@ Abstract.
This study characterizes GPU resilience in Delta  [^1] Delta is an HPC system operated by the National Center for Supercomputing Applications (NCSA) at the University of Illinois Urbana-Champaign., a large-scale AI system that consists of 1,056 A100 and H100 GPUs, with over 1,300 petaflops of peak throughput. We used 2.5 years of operational data (11.7 million GPU hours) on GPU errors. Our major findings include: (i) H100 GPU memory resilience is worse than A100 GPU memory, with $3.2\times$ lower per-GPU MTBE for memory errors, (ii) The GPU memory error-recovery mechanisms on H100 GPUs are insufficient to handle the increased memory capacity, (iii) H100 GPUs demonstrate significantly improved GPU hardware resilience over A100 GPUs with respect to critical hardware components, (iv) GPU errors on both A100 and H100 GPUs frequently result in job failures due to the lack of robust recovery mechanisms at the application level, and (v) We project the impact of GPU node availability on larger-scales and find that significant overprovisioning of 5% is necessary to handle GPU failures.
@@
journalyear: 2025copyright: ccconference: The International Conference for High Performance Computing, Networking, Storage and Analysis; November 16–21, 2025; St Louis, MO, USAbooktitle: The International Conference for High Performance Computing, Networking, Storage and Analysis (SC ’25), November 16–21, 2025, St Louis, MO, USAdoi: 10.1145/3712285.3759821isbn: 979-8-4007-1466-5/2025/11ccs: Computer systems organization Reliabilityccs: Computer systems organization Availabilityccs: Computer systems organization Redundancyccs: Computing methodologies Discrete-event simulationccs: General and reference Measurementccs: General and reference Reliability
@@ 1. Introduction
Large-scale HPC systems are important not only for scientific workloads (skinner2005understanding) but also for data analytics (alam2022survey) and machine learning (ML) (huyen2022designing). The main components of these systems are specialized accelerators, such as GPUs, that enable acceleration of computations, such as ML training (touvron2023llama; team2023gemini), ML inference (li2024llm), and simulations (phillips2005scalable; shainer2009weather).
This paper studies the resilience of A100 GPUs and compares the result with H100 GPUs of the GH200 Grace Hopper Superchips [^2] GH200 Superchip, hereafter., together with their associated memory: 40 GB HBM2e on each A100 GPU, and 96 GB HBM3 on each H100 GPU, respectively. The A100 GPU nodes and H100 GPUs (GH200 Superchip) nodes are operated as two independent systems sharing a storage cluster running the Lustre file system, allowing us to study and compare the two systems (refer to as Delta). The workflow on Delta (delta_2025) involves users from universities nationwide and presents a spectrum of HPC and ML workloads. The study uses 2.5 years of data on critical GPU errors collected across the stated GPUs, encompassing 9.6 million GPU hours of A100 GPUs and 2.1 million GPU hours of H100, a combined 11.7 million GPU hours.
@@
Figure 1. A double-bit memory error (XID 48) occurred, and it is uncorrectable by the SECDED ECC HBM3 memory. Due to this double-bit error, the user job scheduled on that GPU failed, as reflected in the scheduler logs. Subsequently, this uncorrectable memory error requires a node draining and reset to complete the row remapping recovery action. The total recovery process for this incident took 19 hours, during which the node was unavailable for accepting new jobs. This incident shows that a GPU error can lead to user job failures and significantly impact node availability.:
@@
(i) H100 shows $3.2\times$ lower per-GPU mean time between errors (MTBE) compared to A100 for uncorrectable ECC memory errors. The per-GB MTBE of the H100’s HBM3 memory is 24% lower ($\sim\!8.5$M hours) than the A100’s HBM2e memory ($\sim\!11.3$M hours). We conjecture that the reduction in memory resilience stems from H100’s higher memory capacity.
@@
(ii) The GPU memory error-recovery mechanisms on A100 and H100 GPUs (e.g., memory row remapping, error containment) (nv_mem_err) improve GPU memory resilience and reduce service interruption. We observed that these mechanisms mitigate (e.g., using memory row remapping) 92% of uncorrectable ECC memory errors on H100 GPUs. However, the memory error-recovery mechanism is insufficient to handle the increase in memory capacity and the corresponding increase in row remapping events on the H100 GPUs.
@@
(iii) H100 GPUs demonstrate significantly improved hardware resilience over A100 GPUs, with respect to critical components such as GSP [^3] A GPU system processor (GSP) is an onboard co-processor that offloads driver tasks from the CPU for latency and performance improvement., NVLink, and PMU SPI [^4] PMU on an NVIDIA’s GPU regulates the frequency, voltage, and power of the GPU based on various factors such as temperature and power cap. SPI stands for serial peripheral interface, which serves as the communication channel between peripheral hardware., which were major sources of job failures in A100 systems. We attribute this to driver-level enhancements and tighter integration (nv_driver_release_notes; nvidia_grace_hopper), which contribute to improved resilience. Specifically, comparing H100 and A100 GPU hardware, we observed (a) a significant reduction of GSP errors on H100 (only 3 cases in our measurement period) and (b) the elimination of PMU SPI error propagations, which on A100 GPUs can lead to MMU errors 88% of times with 90% leading to user job failures, (c) no NVLink errors on H100 GPUs during the measurement period.
@@
(iv) GPU errors on both A100 and H100 GPUs frequently result in job failures due to the lack of robust recovery mechanisms at the application level. Except for MMU and NVLink errors, other GPU errors cannot be handled by application-level mechanisms, resulting in close to $100\%$ job failure rate. The underlying cause of job failures differs by GPU type: hardware errors are the predominant cause in A100 GPUs, whereas memory errors are the primary cause in H100 GPUs. Overall, we find that application-based recovery strategies are largely ineffective; hence, there is a compelling need to improve resilience at the GPU memory and hardware level.
@@
(v) The overall availability per-GPU node is approximately $99.4\%$ for A100 GPUs and $99.3\%$ for H100 GPUs, corresponding to a downtime between 9–10 minutes per day. We projected the impact of this measured availability on larger scales via emulation. For example, to maintain 99.9% availability at the job level, overprovisioning of 5% would be necessary. While at first glance, such overprovisioning would appear to be a small cost, for the above example, it would cost over \$1 million per month. If GPU node availability were improved to 99.9%, the required overprovisioning would reduce by $2.5\times$.
@@
Putting the paper in perspective. Previous studies on characterizing GPU resilience in large-scale systems (debardeleben2014gpu; di2014lessons; tiwari2015understanding; tiwari2015reliability; gupta2015understanding; gupta2017failures; ostrouchov2020gpu; nie2016large; nie2017characterizing; oles2024understanding) focus on GPU memory errors in older GPU generations (Tesla, Kepler, and Volta) that lack the latest resilience (e.g., row remapping, error containment, NVLink CRC-retry) and performance features (e.g., GPU System Processor) introduced in NVIDIA Ampere-generation GPUs. A recent study from Meta (kokolis2024revisiting) characterizes cluster-level resilience for two large-scale machine learning clusters equipped with A100 GPUs. Our paper provides a deeper understanding of GPU errors and failures of two recent generations of GPUs, A100 and H100, and their impact on a broad set of HPC/ML applications. To the best of our knowledge, this is the first study on A100 and H100 GPU errors in HPC/ML systems.
All software, datasets, and analysis scripts are available in the AD/AE materials and at https://doi.org/10.5281/zenodo.15287639.
@@ 2. Background
This section provides information on (i) Delta specification, (ii) Delta overall GPU utilization, GPU operational environment, and workloads, (iii) critical GPU error categories used in this study, and (iv) GPU error management and recovery.
Figure 2. System architecture and specifications of Delta. This study focuses on the H100 and A100 GPU nodes.:
2.1. Delta Specifications
@@
Note that the H100 GPUs studied here are integrated in GH200 Superchips, tightly coupled to NVIDIA Grace CPUs via NVLink-C2C interconnect, distinguishing them from discrete H100 GPUs. While the H100 microarchitecture and memory specifications are consistent across both variants, differences in CPU–GPU integration may lead to variations in resilience characteristics.
2.2. Delta GPU Operational Environment
A100 and H100 GPUs ran in comparable operational conditions in terms of (i) utilization, (ii) cooling/temperature, and (iii) workloads.
Overall Utilization. Delta’s NVIDIA A100 GPUs are frequently scheduled and utilized, with an average GPU utilization of 51% during the operational period. NVIDIA H100 GPUs show a slightly lower average utilization (41%) than A100 GPUs.
Cooling and Temperature. Delta’s A100 and H100 GPUs are liquid-cooled from shared facility water supply through independent cooling loops. At average utilizations, their mean temperatures are 40C (A100) and 37C (H100), with maximums of 49C and 48C, respectively, indicating comparable operational conditions.
Workloads. We use the allocated Delta projects’ Field-of-Science distributions as a proxy for workload characterization. Both GPUs handled similarly diverse workloads with comparable distributions across the top five fields (A100s/H100s): Computer Science (30.4/32.8%), AI & Intelligent Systems (18.1/29.1%), Applied CS (4.7/4.2%), Biophysics (5.1/1.7%), Materials Engineering (3.6/1.4%) for A100/H100, respectively.
@@ 2.3. NVIDIA GPU Error Categories
NVIDIA GPU errors are reported as XID errors. In this study, we selected a subset of XID errors that are described as common and high-impact by NVIDIA’s Developer Manuals (nv_xid_doc; nv_mem_err), NVIDIA Developer Forums and Blogs, and Delta site reliability engineers (SREs). We primarily collected errors and their associated recovery events. The selected XID errors/events indicate GPU issues that often cannot be resolved without SRE’s interventions (e.g., node service and GPU replacement). The selected XIDs and their corresponding GPU errors are described in Section 4,  Table 1. We categorize the selected GPU errors into three categories: (i) GPU hardware, which includes all onboard hardware except for GPU memory and NVLink interconnect, (ii) NVLink interconnect, and (iii) GPU memory. Note that General GPU Software Error (XID 13) and Reset Channel Verification Error (XID 43) are usually caused by user jobs and do not impact the health of the GPU (nv_xid_doc); we excluded those errors from our study.
GPU Hardware Errors. The critical GPU hardware errors we studied include MMU [^5] The memory management unit (MMU) provides essential memory I/O functionalities. errors, GPU Fallen Off the Bus errors, GSP errors, and PMU communication errors. We do not consider other GPU hardware errors in our study. GPU hardware errors can lead to user job failures, GPU halt, and data corruption. Among those errors, GPU Fallen Off the Bus and GSP errors lead to GPU failures, and manual GPU resets or node reboots are required to recover from the error (nv_xid_doc). Delta SREs monitor GSP errors closely to ensure timely recovery to maintain GPU availability.
GPU Interconnect (NVLink) Errors. GPU-GPU NVLink errors are caused by faulty GPU hardware, connectors, or improper connector installation during system integration, and can lead to GPU unavailability and user job failures. NVLink errors impede data transfer between GPUs and reduce computational throughput. A GPU reset or node reboot is required to clear NVLink errors (nv_xid_doc).
GPU Memory Errors. GPU memory errors included in this study are double-bit errors (DBEs) and consecutive single-bit errors (SBEs) [^6] Consecutive SBEs are multiple single-bit error (SBE) occurrences at the same memory location.. Individual SBEs are not logged, as they are automatically corrected by ECC. DBEs and consecutive SBEs are considered uncorrectable ECC memory errors by the NVIDIA driver, and they trigger downstream error-recovery mechanisms (nv_xid_doc; nv_mem_err), which are introduced in Section 2.3. Failures in these mechanisms can lead to GPU/node failures and require GPU or node reboots to recover (nv_xid_doc). Delta SREs continuously monitor uncorrectable ECC memory errors and error-recovery failures to ensure timely replacement of faulty GPUs.
Figure 3. NVIDIA memory error recovery process for A100 and H100 GPUs.:
@@ 2.4. NVIDIA GPU Error Management
GPU Memory. Figure 3 shows the uncorrectable ECC memory error-recovery process (nv_mem_err) for A100 and H100 in more detail. The primary mechanism for mitigating uncorrectable ECC memory errors for A100 and H100 GPUs is row-remapping, wherein the faulty memory row is replaced with a spare row, and a row-remapping event (RRE) is logged. The actual row remapping happens at the next GPU reset (e.g., during node reboot or maintenance). If there are no spare memory rows, a row remapping failure (RRF) is indicated  (nv_xid_doc; nv_mem_err).
A100 and H100 GPUs support online recovery mechanisms such as error containment and dynamic page offlining (nv_xid_doc; nv_mem_err) for mitigating uncorrectable ECC memory errors with minimal node interruption. The dynamic page offlining marks the faulty memory page as unusable without requiring a GPU reset to maintain availability. The error containment procedure terminates user processes using the faulty memory address to prevent error propagation to other applications. Successful error containment is logged as a Contained Memory Error, whereas an unsuccessful error containment is logged as an Uncontained Memory Error. Failure in a row-remapping or error containment can cause a GPU failure that requires a GPU reset or node reboot. Delta SREs monitor row-remapping failures and replace GPUs that repeatedly emit such errors.
@@ 3.1. Data Sources
Figure 4. Overview of our data collection, processing, and analysis pipeline.:
@@
System logs. System logs collected from all compute nodes capture events across system components. We created a set of regular expression (RegEX) patterns and used it to extract GPU error-recovery log entries by referring to NVIDIA XID messages (nv_xid_doc) from the system logs (Figure 4, item (1)). The GPU error logs were our major sources of error and recovery information.
Slurm scheduler database. Delta uses the Slurm Workload Manager (slurm_log) (“Slurm scheduler” hereafter) for scheduling user jobs. The Slurm scheduler database keeps track of user job scheduling events, including the start and end times, the scheduled nodes, resource usage, job status, exit codes, and the srun command line. We used the Slurm database for user job failure characterization.